Li has two naturally occurring isotopes, 6


Li (isotopic mass 6.0151 amu) and 7


Li (isotopic mass 7.0160 amu).

Calculate the percentage abundance of each isotope if the atomic mass of lithium is 6.941 amu.


1
Expert's answer
2021-08-22T07:30:04-0400

Average atomic mass of an element is defined as the sum of masses of each isotope each multiplied by their natural fractional abundance.


Formula used to calculate average atomic mass follows:



Let the fractional abundance of Li isotope-1 be 'x' and the fractional abundance of Li isotope-2 will be '100-x'


For Li isotope-1 :


Mass of Li isotope-1 = 6.01512 amu


Fractional abundance of Li isotope-1 = x


For Li isotope-2 :


Mass of Li isotope-2 = 7.01600 amu


Fractional abundance of Li isotope-2 = 100-x


Average atomic mass of Li = 6.941 amu


Putting values in equation 1, we get:



By solving the term 'x', we get:



Percent abundance of Li isotope-2 = 100 - x = 100-6.94 = 93.1 %